OpenStudy (anonymous):

Determine whether each curve for y=4x^2 is symmetric with respect to the x-axis, the y-axis and the origin.

OpenStudy (anonymous):

So, what have you tried so far?

OpenStudy (anonymous):

It's actually making alot of sense. So far, all my answers are correct in the back of the book, so now I want to make sure i'm getting the others right. I didn't understand the entire concept until now so its coming together quite well thanks to you. I'm going to submit in here what I have so far, so maybe you could tell me if it looks right if you would be so kind.

OpenStudy (anonymous):

y=4x^2 To solve for the x-axis, I have -y=4x^2 I cannot do anything else with this equation and since it is not the original equation, I am saying there is no x-axis symmetry.

OpenStudy (anonymous):

To solve for y-axis, I have y=(-4x) (-4x) y=4x^2 Original Equation makes it y-axis symmetric

OpenStudy (anonymous):

To solve the origin (replacing with-x, y), I have y=(-4x) (-4x) y=4x^2 Origin YES........ I sure hope i'm doing this a bit better...what do you think?

OpenStudy (anonymous):

yes you right

OpenStudy (anonymous):

Really? That is awesome. I sure have been studying hard considering this isn't even my work. I passed Algebra in 2002 in college with a 100%; however apparently after all those years you forget alot. Its pretty neat though.

OpenStudy (anonymous):

yes its true but when you learned something and forget through some time you remember this very fast

OpenStudy (anonymous):

Im not so sure that I got the origin right though. Aren't I replacing x,y with (-x,-y)

OpenStudy (anonymous):

no let's see for example y=x^3 x-axis is -y=x^3 y-axis is y=(-x)^3

OpenStudy (anonymous):

So original equation is y=4x^2 For the origin, shouldn't I be replacing the x,y with (-x, -y)? Making it: -y=(-4x) (-4x) -y=4x^2 Making the origin not symmetric right?

OpenStudy (anonymous):

what don't you understand

OpenStudy (anonymous):

y-axis for y=4x^2 is origin function

OpenStudy (anonymous):

Okay, so correct the first time. Ugh.

OpenStudy (anonymous):

I put that on my calculator as well and got what you got as well. Thank you very much again.

OpenStudy (anonymous):

Okay, so correct the first time. Ugh.